OpenStudy (anonymous):

please can some one show the working to this problem. The current through a 100-mF capacitor is i(t) = 50sin120pit mA. Calculate the voltage across it at t = 1 ms and t = 5 ms. Take v(0) = 0.

OpenStudy (kenljw):

The voltage across a capacitor is (1/C) Integral of idt

OpenStudy (anonymous):

yes i solved it that way but i keep getting wrong answer. that's why am asking for some 1 to show workings so i can figure out where am wrong. @KenLJW thanks for yor reply.

OpenStudy (kenljw):

1/C = 10 Integral of 50sin(120XpiXt) =50X120XpiXcos(120XpiXt) 0 to 1 ms is 50X120XpiXcos(.12Xpi - 1) same procedure for 5 ms

OpenStudy (anonymous):

@KenLJW thanks alot best response given.

OpenStudy (anonymous):

@KenLJW this is the answer from the text book am using for 1 ms = 93.14 mV, 5ms = 1.736 V. your equation can't reproduce same answer.

OpenStudy (kenljw):

\[1/.1 \int\limits_{0}^{.001} 50\sin(120pit)dt=(500/120\pi)(1-\cos(.12\pi))\] mV = 93.1368 mV, I used wrong Integral at first, when calculating make sure you use radians not degrees.